# Causal Panel Analysis under Parallel Trends:

### v4

@@ Abstract
Two-way fixed effects (TWFE) models are widely used in political science to establish causality, but recent methodological discussions highlight their limitations under heterogeneous treatment effects (HTE) and violations of the parallel trends (PT) assumption. This growing literature has introduced new estimators and diagnostics, causing confusion among researchers about the reliability of existing results and best practices. To address these concerns, we replicated and reanalyzed 49 articles from leading journals using TWFE models for observational panel data with binary treatments. Using six HTE-robust estimators, diagnostic tests, and sensitivity analyses, we find: (i) HTE-robust estimators yield qualitatively similar but highly variable results; (ii) while a few studies show clear signs of PT violations, many lack evidence to support this assumption; and (iii) many studies are underpowered when accounting for HTE and potential PT violations. We emphasize the importance of strong research designs and rigorous validation of key identifying assumptions.
@@
1.  Introduction
Over the past decade, political scientists have increasingly relied on panel data to draw causal conclusions (Xu, 2023). A favored method for such analyses is the two-way fixed effects (TWFE) model because of its ability to control for unobserved time-invariant confounders and common time trends. In our survey of 102 articles published from 2017 to 2023 in three top political science journals using observational panel data with dichotomous treatments, 64 studies (63%) assume a TWFE model with the following functional form or a close variant: [^1] The remaining 38 articles can be categorized into five groups: articles focusing on interaction effects (8 articles), articles using nonlinear links such as logit and Poisson (5 articles), articles employing instrumental variables or regression discontinuity designs (8 articles), and articles using other linear specifications, such as only one-way fixed effects or lagged dependent variables (17 articles). In some articles classified as using TWFE models, “unit” fixed effects are specified at the group level $g$, where multiple units $i$ are nested (e.g., county fixed effects when $i$ indexes cities), or time fixed effects are at a higher level $p$ (e.g., year fixed effects when $t$ indexes days). For simplicity, we use the notation $\alpha_{i}$ and $\xi_{t}$ rather than the more general $\alpha_{g}$ and $\xi_{p}$. For all $i,t$,
$Y_{i,t}=\delta^{TWFE}D_{i,t}+X_{i,t}{{}^{\prime}}\beta+\alpha_{i}+\xi_{t}+\varepsilon_{i,t},$ (1)
where $Y_{i,t}$ and $D_{i,t}$ are the outcome and treatment variables for unit $i$ at time $t$; $X_{i,t}$ is a vector of covariates; $\alpha_{i}$ and $\xi_{t}$ are unit and time fixed effect; and $\varepsilon_{i,t}$ is idiosyncratic errors. Researchers typically interpret $\delta^{TWFE}$ as the treatment effect and estimate the model using ordinary least squares. The resulting estimator for $\delta^{TWFE}$ is commonly known as the TWFE estimator. Moreover, researchers frequently conflate this model with a difference-in-differences (DID) design, and use the two terms interchangeably. [^2] We use the phrase “DID design” in reference to DID research design, which differs from the typical usage in the statistics literature that refers to treatment assignment mechanism.
@@ 1.  Introduction
in which $\mathbf{D_{i}}=\{D_{i,1},D_{i,2},\cdots,D_{i,T}\}$ and $\mathbf{X_{i}}=\{X_{i,1},X_{i,2},\cdots,X_{i,T}\}.$ It means that once current treatment status, covariates, and fixed effects are accounted for, treatment status in any other periods has no partial effect on $Y_{i,t}$ (Wooldridge, 2010, p. 253). Under Model (1), strict exogeneity implies the parallel trends (PT) assumption:
@@
in which $Y_{i,t}(0)=Y_{i,t}(d_{i,t}=0)$ represents the untreated potential outcome for unit $i$ at time $t$. It states that the change in untreated potential outcomes between any two periods is mean independent of the change in observed treatment status during those periods, once changes in covariate values are controlled for. Threats to PT, such as the presence of time-varying confounders and feedback from past outcomes to current treatment assignment, also invalidate strict exogeneity. Therefore, throughout the rest of the paper, we use the term “PT violations” to encompass violations of strict exogeneity. [^3] We discuss the relationship between strict exogeneity and PT, as well as other assumptions, under Model (1) in Section A.1.4 in the Supplementary Materials (SM). Note that the PT assumption invoked by many HTE-robust estimators does not depend on Model (1).
@@
The second group of criticism concerns the consequences of heterogeneous treatment effects (HTE), i.e., $\delta^{TWFE}$ is not a constant (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Athey and Imbens, 2022; Borusyak, Jaravel and Spiess, 2024). Researchers have shown that, under HTE, TWFE estimates in general do not converge to a convex combination of the individual treatment effects (ITE) for units under the treatment condition, even when the PT assumption is valid. The so-called “negative weighting” problem, as described in de Chaisemartin and D’Haultfœuille (2020), is an extremely alarming theoretical result because it implies that a TWFE estimand can be negative (positive) even when all ITE are positive (negative). To address this issue, researchers have proposed many new estimators that are “HTE-robust”—that is, estimators that converge to some convex combinations of ITE under their identifying assumptions.
@@
This paper thus pursues two goals. First, we explain and compare six recent proposals to amend TWFE models, including the interaction weighted (IW) estimator (Sun and Abraham, 2021), stacked DID (Baker, Larcker and Wang, 2022), CS-DID (Callaway and Sant’Anna, 2021), DID multiple (de Chaisemartin and D’Haultfœuille, 2020), PanelMatch (Imai, Kim and Wang 2023, hereafter IKW, 2023), and the imputation method (Borusyak, Jaravel and Spiess 2024, hereafter BJS, 2024; Liu, Wang and Xu 2024, hereafter LWX, 2024). These estimators produce causally interpretable estimates under HTE and PT (or its variants). Second, we replicate and reanalyze 49 studies published in the American Political Science Review (APSR), American Journal of Political Science (AJPS), and The Journal of Politics (JOP) from 2017 to 2023 which rely on a TWFE model to draw causal conclusions. Our aim is to assess the consequences of using or not using HTE-robust estimators and shed light on the prevalence of PT violations in political science research.
@@
Our reanalysis shows that there is considerable variation in the degree to which HTE-robust estimators agree with the TWFE estimator. In most studies, the estimators produce qualitatively similar estimates. A non-negligible minority, however, exhibit significant disagreements. In three cases, at least one of the HTE-robust estimators produces an estimate of the opposite sign as the TWFE estimate; for one of these cases, one estimate of the opposite sign is also statistically significant at 5%. There is also a more widespread problem of power: HTE-robust estimators tend to have larger measures of uncertainty, which, combined with even small fluctuations in point estimates, can weaken statistical confidence. This is especially relevant for results that originally teeter on the brink of significance.
@@
The primary concern, however, is the validity of the PT assumption. While only a few studies show clear signs of PT violations, which likely lead to spurious findings, most studies lack the power to rule out that realistic PT violations could explain a non-zero estimated causal effect. In such cases, even mild PT violations prevent us from concluding that the original treatment effect is nonzero. This does not imply that these studies are incorrect; it simply means that the available data lack sufficient power to reject the null hypothesis of no effects when the PT assumption does not perfectly hold.
@@
Overall, we find that about one-fourth of the articles in our sample meet our standard of being highly credible. In these articles, we can statistically distinguish the treatment effect from zero using an HTE-robust estimator, even allowing for mild PT violations benchmarked against placebo estimates using pre-treatment data. To be clear, we do not intend to criticize the authors of articles we have replicated, as many of the methods we employed were not available when their work was conducted. Our aim is to guide and improve future research.
@@
Most importantly, we urge researchers to prioritize a strong research design and sufficient power in causal panel studies. Credible observational studies should feature a well-defined treatment-outcome pair, shock-induced variation in treatment assignment, and enough power to ensure results are not undermined by small perturbations of key identifying assumptions. Research design has often been overlooked in causal panel analyses, likely because researchers have become accustomed to accepting the strong assumptions behind TWFE models. Recent studies have emphasized the importance of (quasi-)randomness in treatment assignment for the robustness of DID findings (e.g., Roth and Sant’Anna, 2023). We offer recommendations on the choice of estimators and diagnostics in the final section.
@@
This paper makes several contributions. First, we propose a typology of various estimators for causal panel analysis. Our typology is based on the settings in which an estimator can be used and how controls are chosen. We also provide a comprehensive comparison of these estimators and show how several proposals are actually equivalent in some circumstances. We hope this discussion helps researchers deepen their understanding of these estimators. Second, we adapt the robust confidence set (CS) approach for sensitivity analysis proposed by Rambachan and Roth (2023) to the setting of counterfactual estimators. We find it highly useful as it avoids the issue of conditional inference–where hypothesis testing conditional on passing a pretest (e.g., a pretrend test) often results in incorrect size (Roth, 2024). Third, our reanalysis both instills confidence in existing political science research conducted with TWFE models when done properly and raises warnings about potential risks, such as the failure of the PT assumption and a lack of power. Our work also contributes to the ongoing conversation on replication and reproducibility in political science (e.g., Eggers et al., 2015; Lall, 2016; Hainmueller, Mummolo and Xu, 2019; Lal et al., 2024).
@@
Our work is closely related to Baker, Larcker and Wang (2022), who evaluate the credibility of a handful of studies with staggered adoption treatments in the finance and accounting literature. It differs in that: (i) we apply a broader range of estimators and diagnostic tests to a much larger number of empirical applications from a more diverse set of settings—notably, over half of these applications involve treatment reversal, where an additional no-carryover-effect assumption may be violated under TWFE models; (ii) our comprehensive review suggests that while weighting issues under HTE is important, the primary threats to causal inference with panel data, we believe, are PT violations and insufficient power. Our work is also related to McKenzie (2020), Roth et al. (2023), Xu (2023), and Arkhangelsky and Imbens (2024), who review and synthesize the recent methodological advancements of DID in the literature. What sets this paper apart is that the application of these innovations to data, which allows us to assess the relative importance of theoretical criticisms and evaluate the robustness of existing findings.
@@
This research has a few limitations. First, we do not examine methods based on sequential ignorability, an alternative identification framework that assumes no unobserved confounding while allowing for dynamic treatment selection based on observables up to the current time period. Second, our analysis does not encompass studies that use continuous treatments, which is common in political science research. Finally, we note that the methodological literature is still developing, and thus our recommendations should be considered provisional as best practices evolve. Despite these limitations, which we aim to address in subsequent studies, we believe our replication and reanalysis exercise provides valuable insights into some of the widespread and significant challenges in using TWFE models for panel data analysis.
2.  TWFE and Its Pitfalls
@@
In this section, we review the pitfalls of TWFE models identified in the literature. In the classic two-group and two-period case, the TWFE estimator $\hat{\delta}^{TWFE}$ is equivalent to the DID estimator, which consistently estimates the average treatment effect on the treated (ATT) under no anticipation and PT even with HTE. These results do not hold more generally in more complex settings with differential treatment adoption times (known as staggered adoption) or treatment reversal, as we will discuss below.
@@
Our survey of the top journals reveals that the TWFE model under Model (1) is the most commonly adopted approach for estimating causal effects using panel data in political science. Fixed effects (FE) models began their rise to prominence in political science in the early 2000s, and criticism promptly followed. For example, in a debate with Green, Kim and Yoon (2001), Beck and Katz (2001) and King (2001) argue that linear FE models often lead to misleading findings because they throw away valuable information in data, ignore rich temporal dynamics, and are incapable of capturing complex time-varying heterogeneities. Moreover, because both the treatment and outcome variables are often serially correlated in a panel setting, researchers have cautioned against using standard error (SE) estimators suitable for cross-sectional data, such as Huber-White robust SEs (Bertrand, Duflo and Mullainathan, 2004). Scholars have also advocated for bootstrap procedures to better control Type I error rates when the number of clusters (units) is small (Cameron, Gelbach and Miller, 2008).
@@
A second body of research explores the implications of HTE with binary treatments within TWFE models (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Callaway and Sant’Anna, 2021; Borusyak, Jaravel and Spiess, 2024; Athey and Imbens, 2022). Most of this literature assumes staggered adoption, but the insights from that setting are still relevant when there are treatment reversals. In Figure 1, we present two simplified examples with staggered treatment adoption. Figure 1(a) represents outcome trajectories in line with standard TWFE assumptions, which not only include PT but also require that the treatment effect be immediate and unvarying across units and over time. In contrast, the right panel portrays a scenario where PT holds, but the constant treatment effect assumption is not met. Various decompositions by the aforementioned researchers reveal that even under PT, when treatments begin at different times (such as in staggered adoption) and treatment effects evolve over time and vary across units, the TWFE estimand is in general not a convex combination of the ITE for observations subjected to the treatment. The basic intuition behind this theoretical result is that TWFE models use post-treatment data from units who adopt treatment earlier in the panel as controls for those who adopt the treatment later (e.g., Goodman-Bacon, 2021). HTE-robust estimators capitalize on this insight by avoiding these “invalid” comparisons between two treated units in post-treatment periods. In the next section, we present a typology of HTE-robust estimators, along with an introduction to the estimators themselves.
A third limitation of the canonical TWFE model is its presumption of no temporal and spatial interference. In most uses of TWFE models, researchers assume that there are no spatial spillovers and that treatment effects occur contemporaneously, hence no anticipation or carryover effects. No anticipation effects means that future treatments do not affect today’s potential outcomes, while no carryover effects means that today’s treatment does not affect future potential outcomes: For any $i,t$,
@@
in which $\mathbf{d}_{i}=(d_{i,1},d_{i,2},\ldots,d_{i,T})$ is unit $i$’s vector of potential treatment conditions in all periods (from period $1$ to $T$) and $Y_{i,t}(\mathbf{d}_{i})$ is the potential outcome in period $t$ given $\mathbf{d}_{i}$. The TWFE model specified in Model (1) satisfies these two assumptions because $Y_{i,t}(\mathbf{d}_{i})=Y_{i,t}(d_{i,t})$. These assumptions are obviously very strong, but they are rarely questioned or tested in practice (Imai and Kim, 2019; Athey and Imbens, 2022; Wang, 2021). Although some recent methods permit arbitrary carryover effects in staggered adoption settings (Strezhnev, 2018; Callaway and Sant’Anna, 2021), these effects are not distinguishable from contemporaneous effects. [^4] See Section A.4. of the SM in Liu, Wang and Xu (2024) for more details. This limitation becomes more complex when treatment reversal is possible, as demonstrated in Figure 1. In Figure 1(b), data in the left panel are consistent with TWFE assumptions, while the right panel shows deviations from PT, constant treatment effect, and the absence of anticipation or carryover effects. What is concerning is that many real-world data resemble the problematic right figure rather than the ideal left one. Scholars have proposed methods to handle arbitrary carryover effects over a limited number of periods in the general setting (IKW, 2023; LWX, 2024). The challenge of addressing spatial spillover effects without strong structural assumptions still persists (Aronow, Samii and Wang, 2020; Wang, 2021), but its resolution is beyond the scope of this paper.
@@
The finest estimand is the ITE, $\tau_{i,t}=Y_{i,t}(1)-Y_{i,t}(0)$, of which there exists one for each observation $(i,t)$. [^6] This is without loss of generality when feedback and interference are excluded. In staggered DID designs, carryover effects are permissible. When potential outcomes are defined in terms of treatment history, $\tau_{i,t}$ is defined as $Y_{i,t}(1)-Y_{i,t}(\infty)$ where $Y_{i,t}(\infty)$ signifies the untreated potential outcome when unit $i$ never undergoes treatment. Most political science research, however, typically focuses on estimating a single summary statistic. Commonly, this is the ATT, which represents ITE averaged over all observations exposed to the treatment condition. In between these extremes of granularity and coarseness are time-varying dynamic treatment effects (DTE), which are across-unit averages of ITE at each time period relative to treatment adoption (e.g., all observations immediately proceeding treatment adoption). In the staggered adoption setting, we can further subdivide by cohort. We use $\delta_{l}$ ($\delta_{g,l}$) to denote the DTE $l$ periods after treatment adoption (for treatment cohort $g$), with $l=1$ to represent the period immediately after treatment adoption. [^7] Some of the authors we reference denote this first post-treatment period with $l=0$. $\delta_{g,l}$ is also what some authors refer to as a cohort average treatment effect on the treated (CATT) (Strezhnev, 2018; Sun and Abraham, 2021) or group-time average treatment effect (Callaway and Sant’Anna, 2021).
@@
Each of the estimators we discuss can be used to estimate $\delta_{l}$. The outcome model analogous to TWFE for estimating DTE is a lags-and-leads specification. For simplicity, we first describe the staggered setting. Let $K_{i,t}=(t-E_{i,t}+1)$ be the number of periods until (when $K_{i,t}\leq 0$) or since unit $i$’s event time at time $t$ (e.g., $K_{i,t}=1$ if unit $i$ switches into treatment at time $t$). Consider a regression based on the following specification:
@@
$Y_{i,t}=\alpha_{i}+\xi_{t}+X^{\prime}_{i,t}\beta+\sum_{\begin{subarray}{c}l=-a\\ l\neq 0\end{subarray}}^{b}\delta_{l}^{TWFE}\cdot\mathbf{1}\left\{K_{i,t}=l\right\}+\delta_{b+}^{TWFE}\mathbf{1}\left\{K_{i,t}>b\right\}\cdot D_{i,t}+\varepsilon_{i,t},\vspace{-1em}$ (2)

### v5

@@ Abstract
Two-way fixed effects (TWFE) models are widely used in political science to establish causality, but recent methodological discussions highlight their limitations under heterogeneous treatment effects (HTE) and violations of the parallel trends (PT) assumption. This growing literature has introduced numerous new estimators and procedures, causing confusion among researchers about the reliability of existing results and best practices. To address these concerns, we replicated and reanalyzed 49 studies from leading journals using TWFE models for observational panel data with binary treatments. Using six HTE-robust estimators, diagnostic tests, and sensitivity analyses, we find: (i) HTE-robust estimators yield qualitatively similar but highly variable results; (ii) while a few studies show clear signs of PT violations, many lack evidence to support this assumption; and (iii) many studies are underpowered when accounting for HTE and potential PT violations. We emphasize the importance of strong research designs and rigorous validation of key identifying assumptions.
@@
Over the past decade, political scientists have increasingly relied on panel data to draw causal conclusions (Xu, 2023). A favored method for such analyses is the two-way fixed effects (TWFE) model because of its ability to control for unobserved time-invariant confounders and common time trends. In our survey of 102 articles published from 2017 to 2023 in three top political science journals using observational panel data with dichotomous treatments, 64 studies (63%) assume a TWFE model with the following functional form or a close variant: [^1] The remaining 38 studies can be categorized into five groups: studies focusing on interaction effects (8 studies), studies using nonlinear links such as logit and Poisson (5 studies), studies employing instrumental variables or regression discontinuity designs (8 studies), and studies using other linear specifications, such as only one-way fixed effects or lagged dependent variables (17 studies). In some studies classified as using TWFE models, “unit” fixed effects are specified at the group level $g$, where multiple units $i$ are nested (e.g., county fixed effects when $i$ indexes cities), or time fixed effects are at a higher level $p$ (e.g., year fixed effects when $t$ indexes days). For simplicity, we use the notation $\alpha_{i}$ and $\xi_{t}$ rather than the more general $\alpha_{g}$ and $\xi_{p}$.
$Y_{i,t}=\tau^{TWFE}D_{i,t}+X_{i,t}{{}^{\prime}}\beta+\alpha_{i}+\xi_{t}+\varepsilon_{i,t},\quad\text{for all }i,t,$ (1)
where $Y_{i,t}$ and $D_{i,t}$ are the outcome and treatment variables for unit $i$ at time $t$; $X_{i,t}$ is a vector of covariates; $\alpha_{i}$ and $\xi_{t}$ are unit and time fixed effect; and $\varepsilon_{i,t}$ is idiosyncratic errors. Researchers typically interpret $\tau^{TWFE}$ as the treatment effect and estimate the model using ordinary least squares. The resulting estimator for $\tau^{TWFE}$ is commonly known as the TWFE estimator. Moreover, researchers frequently conflate this model with a difference-in-differences (DID) design, and use the two terms interchangeably. [^2] We use the phrase “DID design” in reference to DID research design, which differs from the typical usage in the statistics literature that refers to treatment assignment mechanism (Xu, Zhao and Ding, 2024).
@@
in which $\mathbf{D_{i}}=\{D_{i,1},D_{i,2},\cdots,D_{i,T}\}$ and $\mathbf{X_{i}}=\{X_{i,1},X_{i,2},\cdots,X_{i,T}\}.$ It means that once current treatment status, covariates, and fixed effects are accounted for, treatment status in any other periods has no partial effect on $Y_{i,t}$ (Wooldridge, 2010, p. 253). Under Model (1), strict exogeneity implies a parallel trends (PT) assumption:
@@
in which $Y_{i,t}(0)=Y_{i,t}(d_{i,t}=0)$ represents the untreated potential outcome for unit $i$ at time $t$. It states that the change in untreated potential outcomes between any two periods is mean independent of the change in observed treatment status during those periods, once changes in covariate values are controlled for. Threats to PT, such as the presence of time-varying confounders and feedback from past outcomes to current treatment assignment, also invalidate strict exogeneity. Therefore, throughout the rest of the paper, we use the term “PT violations” to encompass violations of strict exogeneity. [^3] We discuss the relationship between strict exogeneity and PT, as well as other assumptions, under Model (1) in Section A.1 in the Supplementary Materials (SM). Note that the PT assumption invoked by many HTE-robust estimators does not depend on Model (1).
@@
The second group of criticism concerns the consequences of heterogeneous treatment effects (HTE), i.e., $\tau^{TWFE}$ is not a constant (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Sun and Abraham, 2021; Callaway and Sant’Anna, 2021; Athey and Imbens, 2022; Borusyak, Jaravel and Spiess, 2024). Researchers have shown that, under HTE, TWFE estimates in general do not converge to a convex combination of the individual treatment effects (ITE) for units under the treatment condition, even when the PT assumption is valid. The so-called “negative weighting” problem, as described in de Chaisemartin and D’Haultfœuille (2020), is an alarming theoretical result because it implies that a TWFE estimand can be negative (positive) even when all ITE are positive (negative). To address this issue, researchers have proposed many new estimators that are “HTE-robust”—that is, estimators that converge to some convex combinations of ITE under their identifying assumptions.
@@
This paper thus pursues two goals. First, we explain and compare six recent proposals to amend TWFE models, including the interaction weighted (IW) estimator (Sun and Abraham, 2021), stacked DID (Baker, Larcker and Wang, 2022), CS-DID (Callaway and Sant’Anna, 2021), DID multiple (de Chaisemartin and D’Haultfœuille, 2020; De Chaisemartin and d’Haultfoeuille, 2024), PanelMatch (Imai, Kim and Wang 2023, hereafter IKW, 2023), and the imputation method (Borusyak, Jaravel and Spiess 2024, hereafter BJS, 2024; Liu, Wang and Xu 2024, hereafter LWX, 2024). These estimators produce causally interpretable estimates under HTE and PT (or its variants). Second, we replicate and reanalyze 49 studies published in the American Political Science Review (APSR), American Journal of Political Science (AJPS), and The Journal of Politics (JOP) from 2017 to 2023 which rely on a TWFE model to draw causal conclusions. [^4] Replication materials for this paper are available for download at Harvard Dataverse (Chiu et al., 2025). Our aim is to assess the consequences of using or not using HTE-robust estimators and shed light on the prevalence of PT violations in political science research.
@@
Our reanalysis shows that, in most studies, the HTE-robust estimators yield qualitatively similar estimates. However, there is considerable variation in how closely these estimators align with TWFE. In three cases, at least one HTE-robust estimator produces an estimate with an opposite sign to the TWFE estimate; in one of these cases, the opposite-sign estimate is also statistically significant at the 5% level. There is also a more widespread problem of power: HTE-robust estimators tend to have larger measures of uncertainty, which, combined with even small fluctuations in point estimates, can weaken statistical confidence. This is especially relevant for results that originally teeter on the brink of significance.
@@
The primary concern, however, is the validity of the PT assumption. While only a few studies show clear signs of PT violations, which likely lead to spurious findings, most studies lack the power to rule out that realistic PT violations could explain a non-zero estimated causal effect. In such cases, even mild PT violations (informed by pre-treatment estimates) prevent us from concluding that the original treatment effect is nonzero. This does not mean that these studies are wrong; rather, it indicates that the available data do not have sufficient power to reject the null hypothesis of no effects when the PT assumption is not perfectly met.
@@
Overall, we find that about a small minority of the studies in our sample meet our criteria of being highly credible. In these studies, we can statistically distinguish the treatment effect from zero using an HTE-robust estimator, even when allowing for mild PT violations benchmarked against placebo estimates using pre-treatment data. We recognize this as a high standard, as most researchers do not account for the power needed for the sensitivity analysis we perform. To be clear, our intent is not to criticize the authors of the studies we have replicated, since many of the methods we used were not available at the time their studies were conducted. Our goal is to guide and improve future research.
@@
We urge researchers to prioritize a strong research design and sufficient power in causal panel studies. Credible observational studies should feature a well-defined treatment-outcome pair, shock-induced variation in treatment assignment, and sufficient power to ensure results are not undermined by small perturbations of key identifying assumptions. Research design has often been overlooked in causal panel analyses, likely because researchers have become accustomed to accepting the strong assumptions behind TWFE models. Recent studies have emphasized the importance of (quasi-)randomness in treatment assignment for the robustness of DID findings (e.g., Roth and Sant’Anna, 2023).
@@
This paper makes several contributions. First, we propose a typology of various estimators for causal panel analysis. Our typology is based on the settings in which an estimator can be used and how controls are chosen. We also provide a comprehensive comparison of these estimators and show how several proposals are equivalent in some circumstances. We hope this discussion helps researchers deepen their understanding of these estimators. Second, we adapt the robust confidence set (CS) approach for sensitivity analysis proposed by Rambachan and Roth (2023) to the setting of counterfactual estimators. We find it highly useful as it avoids the issue of conditional inference–where hypothesis testing conditional on passing a pretest (e.g., a pretrend test) often results in incorrect size (Roth, 2022). Third, our reanalysis instills confidence in existing political science research that uses TWFE models correctly while also cautioning against potential risks, such as the failure of the PT assumption and insufficient power. Based on these findings, we provide recommendations to improve practices, including the choice of estimators and the use of proper diagnostics. Finally, we contribute to the ongoing conversation on replication and reproducibility in political science (e.g., Eggers et al., 2015; Lall, 2016; Hainmueller, Mummolo and Xu, 2019; Lal et al., 2024).
@@
Our work is closely related to Baker, Larcker and Wang (2022), who evaluate the credibility of a handful of studies with staggered adoption treatments in finance and accounting. It differs in that: (i) we use a wider range of estimators and diagnostic tests on a larger and more diverse set of empirical applications, many of which involve treatment reversals; (ii) our review suggests that while the weighting issue under HTE is important, the main threats to causal inference with panel data are PT violations and insufficient power. Our work also relates to Roth et al. (2023), Xu (2023), de Chaisemartin and D’Haultfœuille (2023), Arkhangelsky and Imbens (2024), and Baker et al. (2025), who review and synthesize the recent methodological advancements in the DID literature. What sets this paper apart is our application of these innovations to data, which allows us to assess the significance of theoretical criticisms.
@@
This research has a few limitations. First, we do not examine methods based on sequential ignorability, an alternative identification framework that assumes no unobserved confounding but allows for dynamic treatment selection up to the current time period. Second, our analysis does not encompass studies that use continuous treatments, which is common in political science research. Finally, as the methodological literature continues to evolve rapidly, our recommendations should be regarded as reflecting current best practices.
TWFE and Its Pitfalls
@@
In this section, we review the pitfalls of TWFE models identified in the literature. In the classic two-group and two-period case, the TWFE estimator $\hat{\tau}^{TWFE}$ is equivalent to the DID estimator, which consistently estimates the average treatment effect on the treated (ATT) under no anticipation and PT even with HTE. These results do not hold more generally in more complex settings with differential treatment adoption times (known as staggered adoption) or treatment reversal, as we will discuss below.
@@
Our survey of the top journals reveals that the TWFE model under Model (1) is the most commonly adopted approach for estimating causal effects using panel data in political science. Fixed effects models began their rise to prominence in political science in the early 2000s, and criticism promptly followed. In a debate with Green, Kim and Yoon (2001), Beck and Katz (2001) and King (2001) argue that linear FE models often lead to misleading findings because they throw away valuable information, ignore rich temporal dynamics, and are incapable of capturing complex time-varying heterogeneities. Moreover, since both treatment and outcome variables are often serially correlated in panel data, researchers have advised against using standard error (SE) estimators suitable for cross-sectional data, such as Huber-White robust SEs (Bertrand, Duflo and Mullainathan, 2004). Scholars also recommend using bootstrap procedures to more effectively control Type I error rates when the number of clusters (units) is small (Cameron, Gelbach and Miller, 2008).
@@
A second body of research explores the implications of HTE with binary treatments within TWFE models (e.g., Goodman-Bacon, 2021; de Chaisemartin and D’Haultfœuille, 2020; Strezhnev, 2018; Callaway and Sant’Anna, 2021; Athey and Imbens, 2022; Borusyak, Jaravel and Spiess, 2024). Most of this literature assumes staggered adoption, but the insights from that setting are still relevant when there are treatment reversals. In Figure 1, we present two simplified examples with staggered treatment adoption. Figure 1(a) represents outcome trajectories in line with standard TWFE assumptions, which not only include PT but also require that the treatment effect be immediate and unvarying across units and over time. In contrast, the right panel portrays a scenario where PT holds, but the constant treatment effect assumption is not met. Various decompositions by the aforementioned researchers reveal that even under PT, when treatments begin at different times (such as in staggered adoption) and treatment effects evolve over time and vary across units, the TWFE estimand is in general not a convex combination of the ITE for observations subjected to the treatment. The basic intuition behind this theoretical result is that TWFE models use post-treatment data from units who adopt treatment earlier in the panel as controls for those who adopt the treatment later (e.g., Goodman-Bacon, 2021). HTE-robust estimators capitalize on this insight by avoiding these “invalid” comparisons between two treated units in post-treatment periods.
A third limitation of the canonical TWFE specification is its presumption of no temporal and spatial interference. In most uses of TWFE models, researchers assume that there are no spatial spillovers and that treatment effects occur contemporaneously, hence no anticipation or carryover effects. No anticipation effects means that future treatments do not affect today’s potential outcomes, while no carryover effects means that today’s treatment does not affect future potential outcomes: For any $i,t$,
@@
in which $\mathbf{d}_{i}=(d_{i,1},d_{i,2},\ldots,d_{i,T})$ is unit $i$’s vector of potential treatment conditions in all periods (from period $1$ to $T$) and $Y_{i,t}(\mathbf{d}_{i})$ is the potential outcome in period $t$ given $\mathbf{d}_{i}$. The TWFE model specified in Model (1) satisfies these two assumptions because $Y_{i,t}(\mathbf{d}_{i})=Y_{i,t}(d_{i,t})$. These assumptions are obviously very strong, but they are rarely questioned or tested in practice (Imai and Kim, 2019; Athey and Imbens, 2022; Wang, 2021). Although some recent methods permit arbitrary carryover effects in staggered adoption settings, these effects are not distinguishable from contemporaneous effects. [^5] See Section A.4 of the SM in Liu, Wang and Xu (2024) for more details. This limitation becomes more complex when treatment reversal is possible, as demonstrated in Figure 1. In Figure 1(b), data in the left panel are consistent with TWFE assumptions, while the right panel illustrates deviations from PT, constant treatment effect, and the absence of anticipation or carryover effects. Real-world data often encounter the challenging scenarios depicted in the right panel rather than the idealized conditions in the left. Scholars have proposed methods to handle limited carryover effects in the general setting (IKW, 2023; LWX, 2024). The challenge of addressing spatial spillover effects without strong structural assumptions still persists (Aronow, Samii and Wang, 2020; Wang, 2021), but its resolution is beyond the scope of this paper.
@@
The finest estimand is the ITE, $\tau_{i,t}=Y_{i,t}(1)-Y_{i,t}(0)$, of which there exists one for each observation $(i,t)$. [^7] This is without loss of generality when feedback and interference are excluded. In staggered DID designs, carryover effects are permissible. When potential outcomes are defined in terms of treatment history, $\tau_{i,t}$ is defined as $Y_{i,t}(1)-Y_{i,t}(\infty)$ where $Y_{i,t}(\infty)$ signifies the untreated potential outcome when unit $i$ never undergoes treatment. Most political science research, however, typically focuses on estimating a single summary statistic. Commonly, this is the ATT, which represents ITE averaged over all observations exposed to the treatment condition. In between these extremes of granularity and coarseness are time-varying dynamic treatment effects (DTE), which are across-unit averages of ITE at each time period relative to treatment adoption (e.g., all observations immediately proceeding treatment adoption). In the staggered adoption setting, we can further subdivide by cohort. We use $\tau_{l}$ ($\tau_{g,l}$) to denote the DTE $l$ periods after treatment adoption (for treatment cohort $g$), with $l=1$ representing the period immediately after treatment adoption. [^8] Some of the authors we reference denote this first post-treatment period with $l=0$. $\tau_{g,l}$ is also what some authors refer to as a cohort average treatment effect on the treated (CATT) (Strezhnev, 2018; Sun and Abraham, 2021) or group-time average treatment effect (Callaway and Sant’Anna, 2021).
@@
Each of the estimators we discuss can be used to estimate $\tau_{l}$. The outcome model analogous to TWFE for estimating DTE is a lags-and-leads specification. For simplicity, we first describe the staggered setting. Let $K_{i,t}=(t-E_{i,t}+1)$ be the number of periods until (when $K_{i,t}\leq 0$) or since unit $i$’s event time at time $t$ (e.g., $K_{i,t}=1$ if unit $i$ switches into treatment at time $t$). Consider a regression based on the following specification:
@@
$Y_{i,t}=\alpha_{i}+\xi_{t}+X^{\prime}_{i,t}\beta+\sum_{\begin{subarray}{c}l=-a\\ l\neq 0\end{subarray}}^{b}\tau_{l}^{TWFE}\cdot\mathbf{1}\left\{K_{i,t}=l\right\}+\tau_{b+}^{TWFE}\mathbf{1}\left\{K_{i,t}>b\right\}\cdot D_{i,t}+\varepsilon_{i,t},\vspace{-1em}$ (2)